In 2020, Dixon Construction Company was awarded a contract to construct an interchange on Highway 30 at a total contract price of $15,000,000. Construction is expected to take two years to complete. The estimated total costs of the project are $12,000,000. Dixon uses the percentage-of-completion method to account for long-term contracts. During the first year cost of construction totaled $7,200,000 and the total estimated costs for the project are unchanged.

a)   Prepare the entry to record 2020 construction costs of $7,200,000 paid in cash.

Dr.  blank      blank  

Cr.  blank      blank

b) Prepare the entry to record the 2020 progress billings of $4,000,000 to the customer.

Dr.  blank      blank  

Cr.  blank      blank

c)   Prepare the entry to record the $3,500,000 payments on the contract received during 2020.

Dr.  blank      blank  

Cr.  blank      blank

d) Prepare the necessary year-end entry to recognize the profit to date.

Dr.  blank      blank
Cr.   blank    blank     

Dr./Cr.  blank      blank

e) What is the amount of gross profit that will appear on the income statement?
blank

Choices:

1,800,000

3,500,000

4,000,000

7,200,000

7,200,000

9,000,000

Cash

Accounts Receivable

Construction Revenue

Billings on contracts

Construction Expense
Answer #1

Solution: (Amount in $)

Part A. Journal Entry to record expenses paid in cash

Date Particular Amount Dr Amount Cr
Construction cost A/c ( Work in Progress ) Dr 7,200,0000
To Cash A/c. 7,200,000

Part B: Progressive billing to customer

Date Particular Amount Dr Amount Cr
Accounts receivable A/c. Dr 4,000,000
To Billings on construction contract 4,000,000

Part C: Payment received

Date Particular Amount Dr Amount Cr
Cash A/c. Dr 3,500,000
To Accounts receivable A/c 3,500,000

Part D:

Date Particular Amount Dr Amount Cr
Profit and Loss A/c Dr 7,200,000
To Construction cost A/c. 7,200,000
( being expenses Transferred to P& L Account )
Construction Revenue A/c. Dr 9,000,000
To Profit And Loss A/c. 9,000,000
( Revenue recognizes based on percentage completion method & transferred to Profit and Loss A/c.)

Note: Net Proft will Be 9,000,000 - 7,200,000 = 1,800,000.

Part E:  

1,800,000 the amount of gross profit that will appear on the income statement

Note:

1. Percentage of completion & revenue to recognize

Percentage of completion = 7,200,000 / 12,000,000 * 100 = 60%

Revenue to recognize

= Percentage of completion * Total revenue of contract

= 60% * 15,000,000 = 9,000,000

2. As there is no indirect expense mentioned in question Net profit and Gross Profit for the year will same.